OpenStudy (anonymous):

How do you get an answer to the equation, |3-X|=20 when solving for X

OpenStudy (anonymous):

set it up like this -20<3-x<20 and you should be able to figure it out

OpenStudy (anonymous):

you can break it into 2 equations: 3 - x = 20 3 - x = -20

OpenStudy (amistre64):

an absolute value doesnt care about the sign of the insides; so it can either be equal to 20, or -20. 3-x = 20 or 3-x = -20

OpenStudy (anonymous):

You have to remember the rule for absolute values: \[|a| = b \implies a = b \text{ or } a = -b\] Then you can split it up into two equations as ninja said and solve each. The union of the two solutions will be all the solutions.

OpenStudy (amistre64):

another option is to take the negative of the innards; which may seem wierd. (3-x) = 20 or -(3-x) = 20